# Exploring Next Token Prediction For Optimizing Databases

### v1

@@ 1. Introduction
Another game-changing factor is the heterogeneity of both the hardware and the workloads. On the hardware front, this heterogeneity manifests in diverse compute and memory chips. For example, modern servers are equipped with large cores, small cores, or a mixture of both. Larger cores are preferable for serialized execution, whereas parallelization suits smaller cores. Likewise, the same Integrated Memory Controllers (IMC) now-a-days support both DRAM and PMEM. In addition, the heterogeneity leads to variable read and write access latencies across hardware, e.g., in Non Uniform Memory Access (NUMA) servers accessing data that resides in a remote DRAM chip can incur up to 3×3$\times$3 × more latency than accessing data that resides in the local DRAM chip (Li et al., 2013). On the workload front, this heterogeneity stems from the diverse range of applications that DBMSs must support, including OLTP, OLAP, HTAP, and modern ML workloads. Each of these hardware and workload applications has distinct performance characteristics, demanding tailored optimization strategies. Adds to this is the resurgence of cloud databases taking over the traditional on-site DBMSs. Amazon Elastic Compute Cloud (EC2) (Amazon, 2024a) provides as many as 750 instance choices equipped with a diverse range of processors, memory, storage, and networks hardware to fit different workload applications.
@@ 2. Database Systems And The Next Token Prediction Paradigm
  • Training-time NTP. During training, the model is fed the previous $\bm{\tau_{<i}}$ ground truth tokens to predict the next token, $\tau_{i}$. The model is provided with the next ground truth token $\tau_{i}$ so that it can maximize the sum of log-probabilites over the entire sequence to better predict $\tau_{i}$ for discrete actions, i.e., $\mathcal{L}_{\text{total}}=-\sum_{i=1}^{T}\log\mathbb{P}(\tau_{i}\mid\bm{\tau_{<i}})$. Notice that, unlike inference, during training the outputs generated by the models are not fed back into itself. As a result, this avoids propagating any compounding error that may result from any misprediction.
@@ 3. The Building Blocks
3.1. Decision Transformer
Decision Transformer (Chen et al., 2021) (DT, for short) is a promising framework that provides the support for modeling sequences conditioned on a particular objective. DT can abstract sequence modeling as a Reinforcement Learning problem (RL, for short), where the sequential decision-making is guided by long-term rewards. It ignores traditional RL techniques, i.e., learning value functions or policies through iterative updates. Rather, it treats RL as a supervised sequence modeling, i.e., an NTP problem. Precisely, a DT is an auto-regressive sequence model that predicts the next action based on the past observations conditioned on a desired reward. By conditioning on a desired reward, DT is able to generate more effective next token predictions, contrary to the traditional NTP paradigm that operates without a specific goal. This reward-conditioned nature of DT allows it to align its next token predictions with optimal decision-making strategies. This aligns well with DBMS tasks with a pre-defined objective, e.g., minimizing query execution time.
@@ 3.1. Decision Transformer
3.2. Tokenization
As discussed in Section 2, the discrete word tokens used in language models do not directly translate to database systems. The word tokens in language models serve as linguistic units that remain consistent across all the language tasks, and datasets. In contrast, database systems lack a universal token representation that makes it harder to adopt the NTP Paradigm.
Ideally, the universal token for database systems (DB-tokens, for short) should be able to generalize across new hardware, and workload applications. Besides, it should be computationally inexpensive to compute so that it does not affect the day-to-day database tasks. Moreover, a DB-token should be able to capture additional context, as unlike language models these action tokens lack any inherent grammar or structure. These requirements present a significant challenge in identifying an effective internal representation for databases. We propose to use as DB-tokens the hardware profiles that are generated from hardware Performance Monitoring Units (PMU, for short). Hardware profiles are computationally inexpensive to retrieve from the hardware registers, and can generalize across different hardware and workload applications, and can provide accurate hardware context that the DBMS is running on at any given time. This makes them an ideal candidate for use as DB-tokens in DBMS optimization tasks, as we demonstrate below.
@@ 3.2. Tokenization
In addition to the DB-tokens, there are action tokens to communicate policies across the DBMS and Machine Learning models. These action tokens represent modular units that make up the policy. The action tokens by themselves hold no inherent meaning. It is the DB-tokens that provide the necessary context, allowing a model to understand and learn the intricacies of DBMS optimization and how specific actions impact performance. In the next section, we describe how to connect these 2 key building blocks, i.e., Decision Transformers and the DB-tokens to develop a framework for adopting the NTP paradigm in database systems.
4. The Framework
@@
  • For a given optimization task with a set objective, i.e., query throughput, tail latency or scalability, we execute different policies across various hardware configurations and workloads, drawn from a diverse set of policy, hardware, and workload pools. This includes both optimal and suboptimal policies.
  • During query execution, we periodically profile the hardware to capture the behavior of important hardware components using hardware statistics, e.g., the number of instructions executed, L1-d, L1-i, Last Level Cache (LLC, for short), and branch misses.
  • Alongside the hardware profiles, we monitor the desired performance metric of the database system. These hardware profiles along with the performance metric serve as the DB-tokens. These DB-tokens are analogous to the state and RTG-tokens in DT. We also tokenize the DBMS policy to generate action tokens.
  • We periodically offload these action tokens along with the associated DB-tokens to an offline dataset for the learning phase.
@@
  • We train a Decision Transformer (DT) on the collected offline dataset in a supervised manner, following the Offline Reinforcement Learning paradigm. As the DT is trained on various policies across varying hardware and workload configurations, it builds a mental model of the optimization task by gaining insights into the inherent tradeoffs and the performance improvements across different strategies, hardware and workload via the interaction of action and DB-tokens.
  • Using the trained Decision Transformer, we infer a new policy auto-regressively. The process begins by feeding DT with the initial DB-token, i.e., state and the desired objective of the specified optimization task. In turn, DT generates the next action token which is fed back to the DBMS to generate the subsequent DB-token. The DB-token is then fed back into DT to get the next action token. This iterative process continues until the complete policy is generated. Note that, during the inference process, the DT has no explicit knowledge about the hardware that the DBMS is running on, or the workload that the DBMS is executing. Hence, both the hardware, and/or the workload can be unseen. It only relies on the DB-tokens, and the inherent knowledge it has gained by being trained on a wide range of policies, hardware configurations, and workloads to generate the next action token.
@@
We formulate main-memory index scheduling as an NTP task. The index is divided into into multiple index chunks, i.e., a set of index nodes. Given a set of compute cores, the goal is to predict the core on which queries should be scheduled for the next index chunk. The data placement is implicit, i.e., the data for each chunk resides in the DIMM chip associated with the assigned core. The ids of the the computing cores serve as the action tokens for scheduling. A complete policy includes a sequence of core ids. The $i$th core in the sequence represents the core assigned to process the corresponding index slice. The clock-cycles, cache, branch misses, local and remote DRAM accesses along with the query throughput constitute the DB-token.
@@ 5. Initial Evaluation
We evaluate the learned scheduling policy by PoLe against various OS policies and heuristics. The default OS policy (OS:D) is local allocation, i.e., the OS kernel places an index node on the NUMA node of the requesting CPU. The OS:I policy places the index nodes in an interleaved manner across all the NUMA nodes. For both the Shared Everything NUMA (SE:N) and the Shared Nothing (Porobic et al., 2012) (SN) strategies, nearby index slices are clustered together and placed in the same NUMA node. For OS:D, OS:I, and SE:N, the OS handles the scheduling. For SN, the scheduling strictly follows the data placement.
We run a YCSB-like benchmark (Cooper et al., 2010) on a main-memory B+-Tree on the YCSB-A workload. Initially, the index is loaded with 30M records. Each index record has a 64-bit key and a 64-bit value. Figure 2 shows the performance of the B+-Tree under different scheduling policies across different hardware. The DT in the above experiment is trained on tokens generated by executing different YCSB workloads, e.g., YCSB-A, YCSB-C, YCSB-E, etc. on a wide range of hardware. Figures 2a, 2b show that PoLe can learn scheduling policies outperforming the baselines by up to 2.78$\times$. Note that, the offline dataset that DT has been trained on contains experiences from different heuristics on the same hardware for the same workload. Thus, this shows PoLe’s ability in learning an optimal policy to improve performance. Figure 2c shows that the scheduling policy learned by PoLe can outperform the competing baselines by up to 3$\times$ on an Intel Skylake X 4 Socket 4 NUMA Node machine. Note that, the offline dataset does not contain any experience from the Intel Skylake X machine. Thus, it shows how the NTP paradigm can help generalize index performance across unseen hardware.

### v2

@@ 1. Introduction
Another game-changing factor is the heterogeneity of both the hardware and the workloads. On the hardware front, this heterogeneity manifests in diverse compute and memory chips. For example, modern servers are equipped with large cores, small cores, or a mixture of both. Larger cores are preferable for serialized execution, whereas parallelization suits smaller cores. Likewise, the same Integrated Memory Controllers (IMC) now-a-days support both DRAM and PMEM. In addition, the heterogeneity leads to variable read and write access latencies across hardware, e.g., in Non Uniform Memory Access (NUMA) servers accessing data that resides in a remote DRAM chip can incur up to 4×4$\times$4 × more latency than accessing data that resides in the local DRAM chip (Bang et al., 2020). On the workload front, this heterogeneity stems from the diverse range of applications that DBMSs must support, including OLTP, OLAP, HTAP, and modern ML workloads. Each of these hardware and workload applications has distinct performance characteristics, demanding tailored optimization strategies. Adds to this is the resurgence of cloud databases taking over the traditional on-site DBMSs. Amazon Elastic Compute Cloud (EC2) (Amazon, 2024a) provides as many as 750 instance choices equipped with a diverse range of processors, memory, storage, and networks hardware to fit different workload applications.
@@ 2. Database Systems And The Next Token Prediction Paradigm
  • Training-time NTP. During training, the model is fed the previous $\bm{\tau_{<i}}$ ground truth tokens to predict the next token, $\tau_{i}$. The model is provided with the next ground truth token $\tau_{i}$ so that it can maximize the sum of log-probabilites over the entire sequence to better predict $\tau_{i}$ for discrete actions, i.e., $\mathcal{L}_{\text{total}}=-\sum_{i=1}^{T}\log\mathbb{P}(\tau_{i}\mid\bm{\tau_{<i}})$. This is termed as Teacher Forcing (Williams and Zipser, 1989). Notice that, unlike inference, during training the outputs generated by the models are not fed back into itself. As a result, this avoids propagating any compounding error that may result from any misprediction.
@@ 3. The Building Blocks
1. Decision Transformer. Decision Transformer (Chen et al., 2021) (DT, for short) is a promising framework that provides the support for modeling sequences conditioned on a particular objective. DT can abstract sequence modeling as a Reinforcement Learning problem (RL), where the sequential decision-making is guided by long-term rewards. It ignores traditional RL techniques, i.e., learning value functions or policies through iterative updates. Rather, it treats RL as a supervised sequence modeling, i.e., an NTP problem. Precisely, a DT is an auto-regressive sequence model that predicts the next action based on the past observations conditioned on a desired reward. By conditioning on a desired reward, DT is able to generate more effective next token predictions, contrary to the traditional NTP paradigm that operates without a specific goal. This reward-conditioned nature of DT allows it to align its next token predictions with optimal decision-making strategies. This aligns well with DBMS tasks with a pre-defined objective, e.g., minimizing query execution time.
@@
2. Tokenization. As discussed in Section 2, the discrete word tokens used in language models do not directly translate to database systems. The word tokens in language models serve as linguistic units that remain consistent across all the language tasks, and datasets. In contrast, database systems lack a universal token representation that makes it harder to adopt the NTP Paradigm.
DB Tokens: Ideally, the universal token for database systems (DB-tokens, for short) should be able to generalize across new hardware, and workload applications. Besides, it should be computationally inexpensive to compute so that it does not affect the day-to-day database tasks. Moreover, a DB-token should be able to capture additional context, as unlike language models these action tokens lack any inherent grammar or structure. These requirements present a significant challenge in identifying an effective internal representation for databases. We propose to use as DB-tokens the hardware profiles that are generated from hardware Performance Monitoring Units (PMU, for short). Hardware profiles are computationally inexpensive to retrieve from the hardware registers, and can generalize across different hardware and workload applications, and can provide accurate hardware context that the DBMS is running on at any given time. This makes them an ideal candidate for use as DB-tokens in DBMS optimization tasks, as we demonstrate below.
@@
Action Tokens: In addition to the DB-tokens, there are action tokens to communicate policies across the DBMS and Machine Learning models. These action tokens represent modular units that make up the policy. The action tokens by themselves hold no inherent meaning. It is the DB-tokens that provide the necessary context, allowing a model to understand and learn the intricacies of DBMS optimization and how specific actions impact performance. In the next section, we describe how to connect these 2 key building blocks, i.e., Decision Transformers and the DB-tokens to develop a framework for adopting the NTP paradigm in database systems.
4. The Probe and Learn (PoLe) Framework
@@
  • Choosing. For a given optimization task with a set objective, i.e., query throughput, tail latency or scalability, we execute different policies across various hardware configurations and workloads, drawn from a diverse set of policy, hardware, and workload pools. This includes both optimal and suboptimal policies.
  • Profiling. During query execution, we periodically profile the hardware to capture the behavior of crucial hardware components using hardware statistics, e.g., the number of instructions executed, L1-d, L1-i, Last Level Cache (LLC, for short), and branch misses.
  • Tokenization. Alongside the hardware profiles, we monitor the desired performance metric of the database system. These hardware profiles along with the performance metric serve as the DB-tokens. These DB-tokens are analogous to the state and RTG-tokens in DT. We also tokenize the DBMS policy to generate action tokens.
  • Offloading. We periodically offload these action tokens along with the associated DB-tokens to an offline dataset for the learning phase.
@@
  • Training. We train a Decision Transformer (DT) on the collected offline dataset following the Teacher Forcing strategy (Williams and Zipser, 1989). As the DT is trained on diverse policies across varying hardware and workload configurations, it builds a mental model of the performance landscape of the optimization task in hand. It gains insights into the inherent tradeoffs and the performance improvements across different strategies, hardware and workload via the interaction of action and DB-tokens, that are present in the offline dataset.
  • Inference. Using the trained DT, we infer a new policy auto-regressively. The process begins by feeding DT with the initial DB-token, i.e., state and the desired objective of the specified optimization task. In turn, DT generates the next action token which is fed back to the DBMS to generate the subsequent DB-token. The DB-token is then fed back into DT to get the next action token. This iterative process continues until the complete policy is generated. Note that, during the inference process, the DT has no explicit knowledge about the hardware that the DBMS is running on, or the workload that the DBMS is executing. Hence, both the hardware, and/or the workload can be unseen. It only relies on the DB-tokens, and the inherent knowledge it has gained by being trained on a wide range of policies, hardware configurations, and workloads to generate the next action token.
@@
We formulate main-memory index scheduling as an NTP task and adopt the PoLe framework to solve it. The index is divided into multiple ($\mathcal{I}_{c}$) index chunks, i.e., a set of index nodes. Given a set of compute cores, the goal is to predict the core on which queries should be scheduled for the next index chunk. The data placement is implicit, i.e., the data for each chunk resides in the DIMM chip associated with the assigned core. The core ids serve as the action tokens. The clock-cycles, cache, branch misses, local and remote DRAM accesses along with the query throughput constitute the DB-token. A complete policy includes a sequence of $\mathcal{I}_{c}$ core ids. The $i$th core in the sequence represents the core assigned to process the corresponding index slice.
@@ 5. Initial Evaluation
We evaluate the learned scheduling policy by PoLe against various OS policies and heuristics. The default OS policy (OS:D) is local allocation, i.e., the OS kernel places an index node on the NUMA node of the requesting CPU. The OS:I policy places the index nodes in an interleaved manner across all the NUMA nodes. For both the Shared Everything NUMA (SE:N) and the Shared Nothing (Porobic et al., 2012) (SN) strategies, nearby index slices are clustered together and placed in the same NUMA node. For OS:D, OS:I, and SE:N, the OS handles the scheduling. For SN, the scheduling strictly follows the data placement. We run a YCSB-like benchmark (Cooper et al., 2010) on a main-memory B+-Tree on the YCSB-A workload. Initially, the index is loaded with 30M records. Each index record has a 64-bit key and a 64-bit value. The DT in the experiment is trained on tokens generated by executing different YCSB workloads, e.g., YCSB-A, YCSB-C, YCSB-E, on a wide range of hardware configuration. Figures 2a, 2b show that PoLe can learn better scheduling policies on Intel Sandy Bridge and NVIDIA Grace Hopper, outperforming the baselines by up to 2.78$\times$. For both these experiments, the workload and the hardware is seen, i.e., the offline dataset that DT has been trained on contains experiences from the same hardware and the same workload. Figure 2c shows that PoLe can learn a better scheduling policy even for an unseen hardware, i.e., the offline dataset that DT is trained on does not contain any experience for the particular hardware. It outperforms the baselines by 3$\times$ on an unseen Intel Skylake X 4 Socket 4 NUMA Node machine. Thus, the NTP paradigm not only enhances performance on existing workloads and hardware but also generalizes to optimize index performance across previously unseen hardware.